# Benchmarking Predictive Coding Networks – Made Simple

### v1

@@ Abstract
In this work, we tackle the problems of efficiency and scalability for predictive coding networks in machine learning. To do so, we first propose a library called PCX, whose focus lies on performance and simplicity, and provides a user-friendly, deep-learning oriented interface. Second, we use PCX to implement a large set of benchmarks for the community to use for their experiments. As most works propose their own tasks and architectures, do not compare one against each other, and focus on small-scale tasks, a simple and fast open-source library adopted by the whole community would address all of these concerns. Third, we perform extensive benchmarks using multiple algorithms, setting new state-of-the-art results in multiple tasks and datasets, as well as highlighting limitations inherent to PC that should be addressed. Thanks to the efficiency of PCX, we are able to analyze larger architectures than commonly used, providing baselines to galvanize community efforts towards one of the main open problems in the field: scalability. The code for PCX is available at https://github.com/liukidar/pcax.
@@ 1 Introduction
The history of predictive coding is long, and spans a large number of disciplines (Friston, 2018; Spratling, 2017). It first appeared as a computational framework in the 50’s, when electronic engineers realized that sending compressed representations of prediction errors in time series data was cheaper than sending the data itself (Elias, 1955). A similar algorithm was then used in the neurosciences, first to describe inhibitory signals in the retina (Srinivasan et al., 1982), and then as a general theory of information processing in different brain regions (Mumford, 1992). In 1999, Rao and Ballard (1999) proposed a hierarchical formulation of predictive coding (PC) as a model of visual processing. Recently, researchers realized that this framework could be used to train neural networks using a bio-plausible learning rule (Whittington and Bogacz, 2017). This has led to different directions of research that either explored interesting properties of PC networks, such as their robustness (Song et al., 2024; Alonso et al., 2022) and flexibility (Salvatori et al., 2022), or proposed variations to improve the performance on specific tasks (Salvatori et al., 2024). While interesting and important for the progress of the field, these lines of research have the tendency of not comparing their results against other papers or those of related fields, and to focus on small-scale experiments. The field is hence avoiding what we believe to be the most important open problem: scaling such results to large scale tasks.
There are multiple reasons why such an important problem has been overlooked. First, it is a hard problem, and it is still unclear why PC is able to perform as well as classical gradient descent with backprop only up to a certain scale, which is of small convolutional models trained to classify the CIFAR10 dataset (Salvatori et al., 2024). Understanding the reason for this limitation would allow us to develop regularization techniques that stabilize learning, and hence allow better performance on more complex tasks, similarly to what dropout and batch normalization have been for deep learning (Srivastava et al., 2014; Ioffe and Szegedy, 2015). Second, the lack of specialized libraries makes PC models extremely slow: a full hyperparameter search on a small convolutional network can take several hours. Third, the lack of a common framework makes reproducibility and iterative contributions impossible, as implementation details or code are rarely provided. In this work, we make first steps toward addressing these problems with three contributions, that we call tool, benchmarking, and analysis.
@@
We release an open-source library for accelerated training for predictive coding called PCX. This library runs in JAX (Bradbury et al., 2018), and offers a user-friendly interface with minimal learning curve through familiar syntax inspired from Pytorch, and extensive tutorials. It is also fully compatible with Equinox (Kidger and Garcia, 2021), a popular deep-learning-oriented extension of JAX, ensuring reliability, extendability, and compatibility with ongoing research developments. It also supports JAX’s Just-In-Time (JIT) compilation, making it efficient and allowing both easy development and execution of PC networks. We empirically show the gain in efficiency with respect to an existing library.
@@
We propose a uniform set of tasks, datasets, metrics, and architectures that should be used as a skeleton to test the performance of future variations of PC. The tasks that we propose are the standard ones in computer vision: image classification and generation (with relative metrics). The models that we use, as well as the datasets, are picked according to two criteria: First, to allow researchers to test their algorithm from the easiest model (feedforward network on MNIST) to more complex ones (deep convolutional models), where we failed to get acceptable results, and should hence pave the way for future research; Second, to favor the comparison against related fields in the literature, such as equilibrium propagation (Scellier and Bengio, 2017). To this end, we have picked some of the models that are consistently used in other research papers. All the source files of the proposed benchmarks, as well as tutorials explaining how to implement them, will be present in the library, to facilitate researchers to use them in their studies.
@@
We provide the baselines for future research by performing an extensive comparative study between different hyperparameters and PC algorithms on multiple tasks. We considered standard PC, incremental PC (Salvatori et al., 2024), PC with Langevin dynamics (Oliviers et al., 2024; Zahid et al., 2023), and nudged PC, as done in the Eqprop literature (Scellier and Bengio, 2017; Scellier et al., 2024). This is also the first time nudging algorithms were applied in PC models. In terms of quantitative contributions, we get state-of-the-art results for PC on multiple benchmarks and show for the first time that it is able to perform well on more complex datasets, such as CIFAR100 and Tiny Imagenet, where we get results comparable to those of backprop. In image generation tasks, we present experiments on datasets of colored images, going beyond MNIST and FashionMNIST as performed in previous works. We conclude with an analysis on the credit assignment of PC, which tries to shed light on problems that we will need to solve to scale up such models even more. To this end, we believe that future effort should aim towards developing algorithms and models that improve the numbers that we show in this work, as they represent the new state of the art in the field.
@@ 2 Related Works
The most related works are those that explore different properties or optimization algorithms of standard PC in the deep learning regime, using formulations inspired by Rao and Ballard’s original work (Rao and Ballard, 1999). Examples are works that study their associative memory capabilities (Salvatori et al., 2021; Yoo and Wood, 2022; Tang et al., 2023, 2024), their ability to train Bayesian networks (Salvatori et al., 2022, 2023b), and theoretical results that explain, or improve, their optimization process (Millidge et al., 2022a, b; Alonso et al., 2022). Results in this field have allowed to either improve the performance of such models in different tasks, or to study different properties that could benefit from the use of PCNs.
@@
In the literature, there are multiple variations of PC algorithms, which differ from Rao and Ballard’s original formulation in the way they update the neural activities. Important examples of such variations are biased competition and divisive input modulation (Spratling, 2008), or the neural generative coding framework (Ororbia and Kifer, 2022). The latter is already used in multiple reinforcement learning and control tasks (Ororbia and Mali, 2023; Ororbia et al., 2023). For a review on how different PC algorithms evolved through time, from signal processing to neuroscience, we refer to (Spratling, 2017); for a more recent review specific to machine learning applications, to (Salvatori et al., 2023a). It is also worth mentioning the original literature on PC in the neurosciences, that does not intersect with ours as it is not related to deep neural networks, has evolved from Rao and Ballard’s work into a general theory that models information processing in the brain using probability and variational inference, called the free energy principle (Friston, 2005; Friston and Kiebel, 2009; Friston, 2010).
@@
Another line of related works is that of neuroscience methods applied to machine learning, like equilibrium propagation (Scellier and Bengio, 2017), which is the most similar to PC (Laborieux and Zenke, 2022; Millidge et al., 2022a). Other methods able to train models of similar sizes are target propagation (Bengio, 2014; Ernoult et al., 2022; Millidge et al., 2022b) and SoftHebb (Moraitis et al., 2022; Journé et al., 2022). The first two communities, that of targetprop and eqprop, consistently use similar architectures in different research papers to test the performance of their methods. In our benchmarking effort, some of the architectures proposed are the same ones they use, to favor a more direct comparison. There are also methods that differ more from PC, such as forward-only methods (Kohan et al., 2023; Nøkland, 2016; Hinton, 2022), methods that back-propagate the errors using a designated set of weights (Lillicrap et al., 2014; Launay et al., 2020), and other Hebbian methods (Moraitis et al., 2022; Journé et al., 2022).
@@ 3 Background and Notation
Predictive coding networks (PCNs) are hierarchical Gaussian generative models with $L$ levels with parameters $\theta=\{\theta_{0},\theta_{1},\theta_{2},...,\theta_{L}\}$, in which each level models a multi-variate distribution parameterized by activation of the preceding level. Let $h_{l}$ be the realization of the vector of random variables $H_{l}$ of level $l$, then we have that the likelihood
@@
For simplicity, we write $P_{\theta_{l}}(h_{l})$ instead of $P_{\theta_{l}}(H_{l}=h_{l})$. We refer to each of the scalar random variables of $H_{l}$ as a neuron. PC assumes that both the prior on $h_{0}$ and the relationships between levels are governed by a normal distribution parameterized as follows:
@@
where $\theta_{l}$ are the learnable weights parametrizing the transformation $f_{l}$, and $\Sigma_{l}$ is a covariance matrix. As it is standard practice, $\Sigma_{l}$ will be fixed to the identity matrix throughout this work (Whittington and Bogacz, 2017). If, for example, $\theta_{l}=(W_{l},b_{l})$ and $f_{l}(h_{l-1},\theta_{l})=\sigma_{l}(W_{l}h_{l-1}+b_{l})$, then the neurons in level $l-1$ are connected to neurons in level $l$ via a linear operation, followed by a non linear map, that is the analogous to a fully connected layer in standard deep learning. Intuitively, $\theta$ is the set of learnable weights of the model, while $h=\{h_{0},h_{1},...,h_{L}\}$ is data-point-dependent latent state, containing the abstract representations for the given observations.
@@
The quantity $\epsilon_{l}=(h_{l}-\mu_{l})$ is often referred to as prediction error of layer $l$, being the difference between the predicted activation $\mu_{l}$ and the current state $h_{l}$. Refer to the appendix, for a full derivation of Eq. (1). To minimize $\mathcal{F}$, the Expectation-Maximization (EM) (Dempster et al., 1977) algorithm is used by iteratively optimizing first the state $h$, and then the weights $\theta$ according to the equations
[…]
We refer to the first step described by Eq. (2) as inference and to the second as learning phase. In practice, we do not train on a single pair $(x,y)$ but on a dataset split in mini-batches that are subsequently used to train the model parameters. Furthermore, both inference and learning are approximated via gradient descent on the variational free energy. In the inference phase, firstly $h$ is initialized to an initial value $h^{(0)}$, and, then, it is optimized for $T$ iterations. Then, during the learning phase we use the newly computed values to perform a single update on the weights $\theta$. The gradients of the variational free energy with respect to both $h$ and $\theta$ are as follows:
@@
Figure 1: Left: Generative and discriminative modes, Right: Pseudocode of one parameter update of PC in supervised learning, as well as an informal description of the different algorithms considered in this work.
@@
This phase is similar to the inference phase, with the difference that we perform it on a test point $\bar{x}$, used to infer the label $\bar{y}$. This is achieved by fixing $h_{0}=\bar{x}$ and compute the most likely value of the latent states $h^{*}\rvert_{h_{0}=\bar{x}}$, again using the state gradients of Eq. (3). We refer to this as discriminative mode. In practice, for discriminiative networks, the values of the latent states computed this way are equivalent to those obtained via a forward pass, that is setting $h_{l}^{(0)}=\mu_{l}^{(0)}$ for every $l\neq 0$, as it corresponds to the global minimum of $\mathcal{F}$ (Frieder and Lukasiewicz, 2022).
@@
So far, we have only described how to use PCNs to perform supervised training. However, as we will see in the experimental section, such models can also be used (and were initially developed to be used) to perform unsupervised learning tasks. Given a datapoint $x$, the goal is to use PCNs to compress the information of $x$ into a latent representation, conceptually similar to how variational autoencoders work (Kingma and Welling, 2013). Such a compression, that should contain all the information needed to generate $x$, is computed by fixing the state vector $h_{L}$ to the data-point, and run inference – that is, we maximize $P_{\theta}(h\rvert_{h_{L}=x})$ via gradient descent on $h$ – until the process has converged. The compressed representation will then be the value of $h_{0}$ at convergence (or, in practice, after $T$ steps). If we are training the model, we then perform a gradient update on the parameters to minimize the variational free energy of Eq.(1), as we do in supervised learning. A sketch of the discriminative and generative ways of training PCNs is represented in Fig. 1(a).
@@ 4 Experiments and Benchmarks
This section is divided in two areas, that correspond to discriminative and generative inference tasks. The first focuses on classification tasks; the second on unsupervised generation. A sketch illustrating the two modes is provided in Fig. 1. To provide a comprehensive evaluation, we will test our models on multiple computer vision datasets, MNIST, FashionMNIST, CIFAR10/100, CelebA, and Tiny ImageNET; on models of increasing complexity, with both feedforward and convolutional layers; and multiple learning algorithms present in the literature.
@@
We consider various learning algorithms present in the literature: (1) Standard PC, already discussed in the background section; (2) Incremental PC (iPC), a simple and recently proposed modification where the weight parameters are updated alongside the latent variables at every time step; (3) Monte Carlo PC (MCPC), obtained by applying unadjusted Langevin dynamics to the inference process; (4) Positive nudging (PN), where the target used is obtained by a small perturbation of the output towards the original, 1-hot label; (5) Negative nudging (NN), where the target is obtained by a small perturbation away from the target, and updating the weights in the opposite direction; (6) Centered nudging (CN), where we alternate epochs of positive and negative nudging. Among these, PC, iPC, and MCPC will be used for the generative mode, and PC, iPC, PN, and NN, and CN for the discriminative mode. See Fig. 1, and the supplementary material, for a more detailed description.

### v2

@@ Abstract
In this work, we tackle the problems of efficiency and scalability for predictive coding networks (PCNs) in machine learning. To do so, we propose a library, called PCX, that focuses on performance and simplicity, and use it to implement a large set of standard benchmarks for the community to use for their experiments. As most works in the field propose their own tasks and architectures, do not compare one against each other, and focus on small-scale tasks, a simple and fast open-source library and a comprehensive set of benchmarks would address all these concerns. Then, we perform extensive tests on such benchmarks using both existing algorithms for PCNs, as well as adaptations of other methods popular in the bio-plausible deep learning community. All this has allowed us to (i) test architectures much larger than commonly used in the literature, on more complex datasets; (ii) reach new state-of-the-art results in all of the tasks and datasets provided; (iii) clearly highlight what the current limitations of PCNs are, allowing us to state important future research directions. With the hope of galvanizing community efforts towards one of the main open problems in the field, scalability, we release code, tests, and benchmarks. [^1] Link to the library: https://github.com/liukidar/pcx
@@ 1 Introduction
In 1999, Rao & Ballard (1999) proposed a formulation of predictive coding (PC) to model hierarchical information processing in the brain. It was recently realized that this framework could be used to train neural networks using a bio-plausible learning rule (Whittington & Bogacz, 2017). This has led to different research directions, whose focus was either to explore interesting properties of PC networks (Song et al., 2024; Alonso et al., 2022), or to propose variations that improve the performance on specific tasks (Salvatori et al., 2024; Ororbia & Kifer, 2022). These lines of research, however, have the tendency of not comparing their results against other works, and to focus on small-scale experiments. The field is hence avoiding what we believe to be the most important open problem: scalability.
There are multiple reasons why the problem of scalability has been overlooked. First, it is a hard problem, and it is still unclear why so far PC has been able to perform as well as classical gradient descent with backpropagation (BP) only up to a certain scale, which is that of small convolutional models trained to classify the CIFAR10 dataset (Salvatori et al., 2024). Understanding this would allow us to develop regularization techniques that stabilize learning, and hence allow better performance on more complex tasks. Second, the lack of specialized libraries makes PC models extremely slow: a full hyperparameter search on a small convolutional network can take several hours. Third, the lack of a common framework makes reproducibility and iterative contributions hard, as implementation details or code are rarely provided. In this work, we make the first steps toward addressing these problems with three contributions, that we call tool, benchmarking, and analysis.
@@
We release an open-source library for accelerated training for predictive coding called PCX. This library runs in JAX (Bradbury et al., 2018), and offers a user-friendly interface with a minimal learning curve through familiar syntax inspired by Pytorch. We also provide extensive tutorials. It is also fully compatible with Equinox (Kidger & Garcia, 2021), a popular deep-learning-oriented extension of JAX, ensuring reliability, extendability, and compatibility with ongoing research developments. It also supports JAX’s Just-In-Time (JIT) compilation, making it efficient and allowing both easy development and execution of PC networks, gaining efficiency with respect to existing libraries.
@@
We propose a uniform set of tasks, datasets, metrics, and architectures that should be used as a skeleton to test the performance of future variations of PC. The tasks that we propose are the standard ones in computer vision: image classification and generation. The models that we use, as well as the datasets, are picked according to two criteria: First, to allow researchers to test their algorithm from the easiest task (feedforward network on MNIST) to more complex ones; Second, to favor the comparison against related fields in the literature, such as equilibrium and target propagation (Scellier & Bengio, 2017; Bengio, 2014). To this end, we have picked some of the models that are consistently used in their research papers. As learning algorithms, we consider standard PC, incremental PC (Salvatori et al., 2024), PC with Langevin dynamics (Oliviers et al., 2024), and nudged PC, as done in the Eqprop literature (Scellier & Bengio, 2017; Scellier et al., 2024). Note that this is the first time nudging algorithms are applied in PC models.
@@
We get state-of-the-art (SOTA) results for PC on multiple benchmarks and show for the first time that it is able to perform well on more complex datasets, such as CIFAR100 and Tiny Imagenet, where we get results comparable to those of backprop. In image generation tasks, we present experiments on datasets of colored images, going beyond MNIST and FashionMNIST as performed in previous works. We thoroughly discuss the results and highlight areas of improvement, the main one being generalization to very deep models, and report analysis on the credit assignment of PC in such cases, to better understand the reasons behind some failures. To conclude, in the supplementary material we provide a detailed explanation of hyperparameters/techniques/tricks that allowed us to reach SOTA results, to also provide a cookbook for researchers in the field.
@@ 2 Related Works
The most related works are those that explore different properties or optimization algorithms of standard PC in the deep learning regime, using formulations inspired by Rao and Ballard’s original work (Rao & Ballard, 1999). Examples are works that study their associative memory capabilities (Salvatori et al., 2021; Yoo & Wood, 2022; Tang et al., 2023; 2024), their ability to train Bayesian networks (Salvatori et al., 2022; 2023b), and theoretical results that explain, or improve, their optimization process (Millidge et al., 2022a; b; Alonso et al., 2022). Results in this field have allowed either to improve the performance of such models in different tasks, or to study different properties that could benefit from the use of PCNs.
@@
In the literature, there are multiple variations of PC algorithms. Important examples are biased competition and divisive input modulation (Spratling, 2008), or the neural generative coding framework (Ororbia & Kifer, 2022). The latter is already used in multiple reinforcement learning and control tasks (Ororbia & Mali, 2023; Ororbia et al., 2023), and has its own JAX-based open source library called NGCLearn. For a review on how different PC algorithms evolved through time, from signal processing to neuroscience, we refer to (Spratling, 2017); for a more recent review specific to machine learning applications, to (Salvatori et al., 2023a). It is also worth mentioning the original literature on PC in the neurosciences has evolved from Rao and Ballard’s work into a general theory that models information processing in the brain using probability and variational inference, called the free energy principle (Friston, 2005; Friston & Kiebel, 2009; Friston, 2010).
@@
Another line of related works is that of neuroscience methods applied to machine learning, like equilibrium propagation (Scellier & Bengio, 2017), which is the most similar to PC (Laborieux & Zenke, 2022; Millidge et al., 2022a). Other methods able to train models of similar sizes are target propagation (Bengio, 2014; Ernoult et al., 2022; Millidge et al., 2022b) and SoftHebb (Moraitis et al., 2022; Journé et al., 2022). The first two communities, that of targetprop and eqprop, consistently use similar architectures in their research papers to test their methods. In our benchmarking effort, some of the architectures proposed are the same ones, to favor a more direct comparison. There are also methods that differ more from PC, such as forward-only methods (Kohan et al., 2023; Nøkland, 2016; Hinton, 2022), and methods that back-propagate the errors using a designated set of weights (Lillicrap et al., 2014; Launay et al., 2020).
@@ 3 Background and Notation
Predictive coding networks (PCNs) are hierarchical Gaussian generative models that consist of $L$ levels. Each level models a multi-variate distribution, parameterized by the activation of the preceding level, which depends on both the model parameters $\theta={\theta_{0},\theta_{1},\theta_{2},...,\theta_{L}}$ and the model state $h$. Let $h_{l}\in h$ be the realization of the vector of random variables $H_{l}$ of level $l$, then we have that the likelihood
@@
Where we write $P_{\theta_{l}}(h_{l})$ instead of $P_{\theta_{l}}(H_{l}=h_{l})$, that is the likelihood of $H_{l}$ evaluated at $h_{l}$. We refer to each of the scalar random variables of $H_{l}$ as a neuron. In PC both the prior on $h_{0}$ and the relationships between levels are governed by a normal distribution parameterized as follows:
@@
where $\theta_{l}$ are the learnable weights parametrizing the transformation $f_{l}$, and $\Sigma_{l}$ is a covariance matrix, that will be fixed to the identity matrix throughout this work. If, for example, $\theta_{l}=(W_{l},b_{l})$ and $f_{l}(h_{l-1},\theta_{l})=\sigma_{l}(W_{l}h_{l-1}+b_{l})$, then the neurons in level $l-1$ are connected to neurons in level $l$ via a linear operation, followed by a non-linear map, analogously to a fully connected layer. Intuitively, $\theta$ is the set of learnable weights of the model, while $h=\{h_{0},h_{1},...,h_{L}\}$ is data-point-dependent latent state, containing the abstract representations for the given observations.
@@
The quantity $\epsilon_{l}=(h_{l}-\mu_{l})$ is often referred to as prediction error of layer $l$, being the difference between the predicted activation $\mu_{l}$ and the current state $h_{l}$. For a full derivation of Eq. (1) we refer to the appendix. To minimize $\mathcal{F}$, the Expectation-Maximization (EM) (Dempster et al., 1977) algorithm is used by iteratively optimizing first the state $h$, and then the weights $\theta$ according to the equations
$h^{*}=\text{argmin}_{h}\mathcal{F}(h,\theta),\;\;\theta^{*}=\text{argmin}_{\theta}\mathcal{F}(h^{*},\theta).$ (2)
We refer to the first step described by Eq. (2) as inference and to the second as learning phase. In practice, we do not train on a single pair $(x,y)$ but on a dataset split into mini-batches that are subsequently used to train the model parameters. Furthermore, both inference and learning are approximated via gradient descent on the variational free energy. In the inference phase, firstly $h$ is initialized to an initial value $h^{(0)}$, and then, it is optimized for $T$ iterations. Then, during the learning phase, we use the newly computed values to perform a single update on the weights $\theta$. The gradients of the variational free energy with respect to both $h$ and $\theta$ are as follows:
@@
Figure 1: (a): Generative and discriminative modes; (b): Pseudocode of PC in supervised learning, where both the latent variables $h_{l}$ and the weight parameters $\theta_{l}$ are updated to minimize the variational free energy $\mathcal{F}$. In the colored boxes, informal description of the different algorithms considered in this work.
@@
Given a test point $\bar{x}$, we fix $h_{0}=\bar{x}$ and compute the most likely value of the latent states $h^{*}\rvert_{h_{0}=\bar{x}}$, again using the state gradients of Eq. (3). We refer to this as discriminative mode. In practice, for discriminative networks, the values of the latent states computed this way are equivalent to those obtained via a forward pass, that is setting $h_{l}^{(0)}=\mu_{l}^{(0)}$ for every $l\neq 0$, as it corresponds to the global minimum of $\mathcal{F}$ (Frieder & Lukasiewicz, 2022).
@@
PCNs can also be used to perform unsupervised learning tasks. Given a data point $x$, the goal is to compress the information of $x$ into a latent representation, conceptually similar to how variational autoencoders work (Kingma & Welling, 2013). Such a compression is computed by fixing the state vector $h_{L}$ to the data point, and running inference – that is, we maximize $P_{\theta}(h\rvert_{h_{L}=x})$ via gradient descent on $h$. The compressed representation will then be the value of $h_{0}$ at convergence (or, in practice, after $T$ steps). If we are training the model, we then perform a gradient update on the parameters to minimize the variational free energy of Eq. (1), as we do in supervised learning. A sketch of the discriminative and generative ways of training PCNs is represented in Fig. 1(a).
@@ 4 Experiments and Benchmarks
The benchmark that we propose is a standardized set of models, datasets, and testing procedures that have been consistently used to evaluate predictive coding, but in a non-uniform way. Here, for a comprehensive evaluation, we test models of increasing complexity on multiple computer vision datasets, with both feedforward and convolutional/de-convolutional layers; and multiple learning algorithms present in the literature. This section is divided into two areas that correspond to discriminative (supervised) and generative (unsupervised) inference tasks. For the former mode, we focus on supervised classification, and unsupervised generation for the latter. A sketch illustrating the two modes is in Fig. 1. For every class of experiments, we have performed a large hyperparameter search, and the details needed to reproduce the experiments, as well as a discussion about ‘lessons learned’ during such a large search, are in the Appendix B and C.
To provide a comprehensive evaluation, we have tested on multiple computer vision datasets, MNIST (LeCun & Cortes, 2010), FashionMNIST (Xiao et al., 2017), CIFAR10/100 (Krizhevsky et al., 2009), CelebA (Liu et al., 2018), and Tiny ImageNET (Le & Yang, 2015); on models of increasing complexity, and multiple learning algorithms present in the literature. The results, averaged over $5$ seeds are reported in Tab. 1 when we used discriminative models, and in Tab. 2 for generative models. Note that, besides a very recent exception on CelebA (Sennesh et al., 2024), this is the first time that PCNs with local message passing are tested on datasets such as CelebA, CIFAR100, and Tiny ImageNet.
@@
We consider various learning algorithms present in the literature: (1) Standard PC, already discussed in the background section; (2) Incremental PC (iPC) (Salvatori et al., 2024), a simple and recently proposed modification where the weight parameters are updated alongside the latent variables at every time step; (3) Monte Carlo PC (MCPC) (Oliviers et al., 2024), obtained by applying unadjusted Langevin dynamics to the inference process; (4) Positive nudging (PN), where the target used is obtained by a small perturbation of the output towards the original, 1-hot label; (5) Negative nudging (NN), where the target is obtained by a small perturbation away from the target, and updating the weights in the opposite direction; (6) Centered nudging (CN), where we alternate epochs of positive and negative nudging (Scellier et al., 2024). Among these, PC, iPC, and MCPC will be used for the generative mode, and PC, iPC, PN, NN, and CN for the discriminative mode. See Fig. 1, and the supplementary material, for a more detailed description.